Part A: supernormal-growth stock valuation

A firm’s cash dividend is expected to grow at the following rates for the next 5 years. From year 6 on, its growth rate stabilizes at 5% into the foreseeable future. The firm’s required rate of return is 11.75%, and its most-recent dividend was at $1.75 per share.

Year

1

2

3

4

5

6 to infinity

Growth rate per year, %

30

25

20

15

10

5

i. Estimate the firm’s current stock price in $wx.yz format

ii. If the stock is trading at $50.00 per share, what is the implied required rate of return? Give answer in % to 4 decimal places. [Hint: use Data, What-if analysis, and Goal seek functions.]

May use IRR
Answer #1

i.

Year Cash Flow Discounting Factor
[1/(1.1175^year)]
PV of Cash Flow
(cash flow*discounting factor)
1 1.75 + 30% = 2.275 0.894854586 2.035794183
2 2.275 + 25% = 2.84375 0.80076473 2.277174702
3 2.8438 + 20% = 3.4125 0.716567991 2.44528827
4 3.4125 + 15% = 3.924375 0.641224153 2.516404037
5 3.9244 + 10% = 4.316813 0.573802374 2.476997262
5 Terminal Value=
(4.3168+5%)/ (11.75%-5%)
67.15022 0.573802374 38.53095695
Expected Share Price today
=sum of PVs
50.2826154

Current Stock Price = $50.28

ii.

10% 11% 12.00%
Period Cash Flow Discountig Factor
[1/(1.1^period)]
PV of cash flows
(cash flow*discounting factor)
Discountig Factor
[1/(1.11^period)]
PV of cash flows
(cash flow*discounting factor)
Discountig Factor
[1/(1.12^period)]
PV of cash flows
(cash flow*discounting factor)
0 -50 1 -50 1 -50 1 -50
1 2.275 0.9090909 2.06818182 0.9009009 2.04954955 0.8928571 2.03125
2 2.84375 0.8264463 2.35020661 0.8116224 2.308051295 0.7971939 2.2670201
3 3.4125 0.7513148 2.56386176 0.7311914 2.495190589 0.7117802 2.4289501
4 3.924375 0.6830135 2.68040093 0.658731 2.585107367 0.6355181 2.4940113
5 4.316813 0.6209213 2.68040124 0.5934513 2.561818408 0.5674269 2.4494756
5 67.15022 0.6209213 41.6950034 0.5934513 39.85038724 0.5674269 38.102838
PV = 4.0380558 PV = 1.850104446 PV = -0.226455

IRR is the rate of return at which NPV=0

Here, NPV@11% is positive and @12% is negative.

Therefore, IRR is between 11% and 12%

IRR = Rate at which positive NPV + [Positive NPV/(Positive NPV-Negative NPV)]

= 11% + [1.8501/(1.8501-(-0.2265)]

= 11% + [1.8501/2.0766]

= 11% + 0.8909% = 11.8909%

(Explanation & Logic of the method: NPV @11% is 1.8501 and NPV@12% is -0.2265. i.e. 1% increase in required rate of return reduces NPV by 1.8501+0.2265 =2.0766. We want NPV=0. Therefore, Proportionate increase in required rate of return to reduce NPV by 1.8501 is calculated)